If 4a+3b=27 and 3a+4b=29, then 7a+7b/2=?
Answer by RAY100(1637)

(1) 4a +3b =27
(2) 3a +4b = 29
.
multiply (1) by 3,,,,and (2) by 4
.
(3) 12a +9b = 81
(4) 12a +16b = 116
.
(4) - (3)
7b = 35
b=5
.
subst into (1)
4a +3(5) = 27
4a =12
a=3
.
check
(1) 4(3) +3(5) = 27,,,,,12 +15 =27 ,,,,,,ok
(2) 3(3) +4(5) =29,,,,9+20=29,,,,,,,ok
.
subst into 7a + 7b/2 = 7(3) +7(5) /2 = 21 + 35/2 = 38.5
.
or subst (7a +7b )/2 = (7*3 +7*5)/2= 56 /2 =28